Sin-1(Sin2)

Sin-1(Sin2)

Grade:12

1 Answers

Chetan Mandayam Nayakar
13 years ago

the angle measuring 2 radians lies between π/2 and 'π' Now the range of the inverse circular sine function is from -π/2 to +π/2

in addition sin(2) = sin(π -2). Therefore sin-1(sin2) = π -2  radians

'Win exciting gifts by answering the questions on Discussion Forum.